OpenStudy (anonymous):

Mary is 24 years old. Mary is twice as old as Ann was when Mary was as old as Ann is now. How old is Ann? a. 6 b. 9 c. 12 d. 15 e. 18

OpenStudy (followthefez):

e) 18

OpenStudy (texaschic101):

this is confusing :/

OpenStudy (anonymous):

why is it 18? D:

OpenStudy (texaschic101):

how did you get 18 ?

OpenStudy (texaschic101):

wouldn't it be 12 ?

OpenStudy (followthefez):

Hold up. Just trying to figure out how to explain!

OpenStudy (texaschic101):

if Ann is 12 now, then mary is twice as old, which is 24. Wouldn't it be 12 ?

OpenStudy (anonymous):

"Mary is twice as old as Ann was" means that she's above 12, right? but thats all i know

OpenStudy (texaschic101):

oh......I see what you are saying 430on Sat....

OpenStudy (anonymous):

hehe :(

OpenStudy (followthefez):

Mary is twice as old as Ann WAS, when Mary WAS as old as Ann is now. Mary IS NOT twice as old as Ann. She is twice as old as Ann was, when SHE was Ann's current age. You have to look at the difference in their ages. The unknown is Ann's age. Let this be y. The difference in their ages is 24-y. So when Mary was x years old, Ann's age was y-(24-y). And Mary is twice the age of Ann, when Ann was y-(24-y). So if you put that all together, Ann's age is 24= 2 x y-(24-y) Solving for y gives 18. (Which also shows that y-(24-y) is 12) Does that make sense?

OpenStudy (anonymous):

wait, i thought she was twice as old as ann was when she was ann's age. D:

OpenStudy (texaschic101):

@jim_thompson5910 ...can you confirm this

OpenStudy (followthefez):

She is. When she was Ann's CURRENT age (18), Ann was 12. That means she is 2x12 years old - 24.

OpenStudy (followthefez):

Let Ann's age equal y The difference in their ages is 24-y <-- let's call this d. So when Mary was Ann's current age, Ann's age was HER current age, minus the difference in their ages. ie y-d And Mary is currently twice the age of Ann, when Ann was (y-d) years old. So if you put that all together, Ann's age is 24= 2 x (y-d) Solving for y gives 18. Does that look cleaner?

jimthompson5910 (jim_thompson5910):

Let x = ann's current age and y = # of years ago where Mary was ann's current age x so if Mary is 24 years now, y years ago Mary would be 24 - y years old (ann's current age), so 24 - y = x x = 24 - y Mary's current age of 24 is twice that of Anns in the past, which is x - y years old so this means 24 = 2*(x-y) 24 = 2*(24 - y-y) ... plug in x = 24 - y 24 = 48 - 4y 24 - 48 = -4y -24 = -4y -4y = -24 y = -24/(-4) y = 6 So this means that 6 years ago, Marry was twice as old as Ann was. So if Mary is 24 now, then Ann must be 24 - 6 = 18 years old This problem is definitely a pain since there's a lot to keep track of and you can easily get lost

OpenStudy (texaschic101):

Fez got it correct.....good job Fez

OpenStudy (texaschic101):

thank you jim

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (followthefez):

Haha. Thanks. Not the easiest thing in the world to solve!